Moving to another question will save this response. «. Questio Question 2 4 points Calculate the freezing point of a solution of 620.0 g of ethylene glycol (C2H602) dissolved in 500.0 g of water. Kf= 1.86emandRD 0.512 C/m. 30 degrees Celsius C-30 degrees Celsius 37.2 degrees Celsius -37.2 degrees Celsius 8 6 7 8 9 0 5
Answer #1

Answer : - 37.2 oC

mass of ethylene glycol = 620.0 g

moles of ethylene glycol = 620 / 62.07 = 9.99

mass of water = 500 g = 0.5 kg

molality = moles / mass of solvent

            = 9.99 / 0.5

molality = 19.98 m

delta Tf = Kf x m

0 - Tf = 1.86 x 19.98

Tf = - 37.2

freezing point = - 37.2 oC